Problem 18-2, p. 1042

P18-2 (LO2,3,4) (Allocate Transaction Price, Modification of Contract) Refer to the Tablet Bundle A revenue arrangement in P18-1 (See below). In response to competitive pressure for Internet access for Tablet Bundle A, after 2 years of the 3-year contract, Tablet Tailors offers a modified contract and extension incentive. The extended contract services are similar to those provided in the first 2 years of the contract. Signing the extension and paying $90 (which equals the standalone selling of the revised Internet service package) extends access for 2 more years of Internet connection. Forty Tablet Bundle A customers sign up for this offer.

Instructions

(a)Prepare the journal entries when the contract is signed on January 2, 2019, for the 40 extended contracts. Assume the modification does not result in a separate performance obligation.

(b)Prepare the journal entries on December 31, 2019, for the 40 extended contracts (the first year of the revised 3-year contract).

Reference Problem:

P18-1 (LO2,3) (Allocate Transaction Price, Upfront Fees) Tablet Tailors sells tablet PCs combined with Internet service, which permits the tablet to connect to the Internet anywhere and set up a Wi-Fi hot spot. It offers two bundles with the following terms.

1.Tablet Bundle A sells a tablet with 3 years of Internet service. The price for the tablet and a 3-year Internet connection service contract is $500. The standalone selling price of the tablet is $250 (the cost to Tablet Tailors is $175). Tablet Tailors sells the Internet access service independently for an upfront payment of $300. On January 2, 2017, Tablet Tailors signed 100 contracts, receiving a total of $50,000 in cash
Answer #1

a.

Date Particulars Debit Credit
2-Jan-19 Cash 3600
Uneraned Service Revenue 3600

Explanation: 3600 = 40*90

Please note that services in the extended period are the same as the services that were provided in the original contract period. As they are not distinct hence the modifications will be considered as part of the original contract.

b.

Date Particulars Debit Credit
31-Dec-19 Uneraned Service Revenue 2413
Service revenue 2413

Explanation: Original internet service contract = 40*273 = 10,290

Revenue recognized in 1st two years = 10290*2/3 = 7280

Remianining service at orignal rates = 10290-7280 = 3640

Extended service = 3600

3640+3600 = $7240

7240/3 = $2413